Dexcon Technologies, Inc., is evaluating two alternatives to produce its new plastic filament with tribological (i.e., low friction) properties for creating custom bearings for 3-D printers. The estimates associated with each alternative are shown below. Using a MARR of 16% per year, which alternative has the lower present worth?

Method DDM LS
First Cost $-170,000 $-350,000
M&O Cost, per Year $-65,000 $-40,000
Salvage Value $4,000 $29,000
Life 2 years 4 years

The present worth for the DDM method is $  .

The present worth for the LS method is $  .
Answer #1


Question

The life of DDM method is 2 years.

The life of LS option is 4 years.

Since, the life of two options is not same, we have to use LCM method to calculate the present worth.

LCM of 2 and 4 is 4.

This means DDM method would be replaced by similar method after two years.

Calculate the present worth of DDM method -

PW = -170,000 - 65,000(P/A, 16%, 4) - 166,000(P/F, 16%, 2) + 4,000(P/F, 16%, 4)

PW = -170,000 - (65,000 * 2.7982) - (166,000 * 0.7432) + (4,000 * 0.5523)

PW = -170,000 - 181,883 - 123,371.2 + 2,209.2

PW = -473,045

The Present Worth of DDM method is -$473,045

Calculate the present worth of LS method -

PW = -350,000 - 40,000(P/A, 16%, 4) + 29,000(P/F, 16%, 4)

PW = -350,000 - (40,000 * 2.7982) + (29,000 * 0.5523)

PW = -350,000 - 111,928 + 16,016.7

PW = -445,911.3

The Present Worth of LS method is -$445,911.3

So,

The Present worth of LS method is lower and thus, LS method should be selected.